OpenStudy (lokiando):

Which polynomial is prime? a) 7x^2 – 35x + 2x – 10 b) 9x^3 + 11x^2 + 3x – 33 c) 10x^3 – 15x^2 + 8x – 12 d) 12x^4 + 42x^2 + 4x^2 + 14

OpenStudy (mathstudent55):

Try factoring by grouping.

OpenStudy (lokiando):

how?

OpenStudy (lokiando):

just give me the answer

OpenStudy (mathstudent55):

Here is a) as an example. Group the first two terms, and group the last two terms. Then try to factor a common factor out of the first group and a common factor out of the second group. 7x^2 – 35x + 2x – 10 = (7x^2 - 35x) + (2x - 10) = 7x(x - 5) + 2(x - 5) = (x - 5)(7x + 2)

OpenStudy (mathstudent55):

Now try the same with the others.

OpenStudy (lokiando):

tnx